Nine uncharged capacitors with equal capacitances are combined in series. The combination is connected to a 5.39 V battery, which charges the capacitors. The charging process involves 0.000291 C of charge moving through the battery. Find the capacitance of each capacitor.
Answer #1

The equivalent capacitance of this circuit which will be given by -

Ceq = Q / V

where, Q = charge moving through a battery = 0.000291 C

V = voltage of a battery = 5.39 V

then, we get

Ceq = [(0.000291 C) / (5.39 V)]

Ceq = 5.39 x 10-5 F

We know that, nine uncharged capacitors with equal capacitances are combined in series.

Then, the capacitance of each capacitor which will be given as -

1 / Ceq = 1 / C1 + 1 / C2 + 1 / C3 + 1 / C4 + 1 / C5 + 1 / C6 + 1 / C7 + 1 / C8 + 1 / C9

1 / Ceq = 9 / C

C = 9 Ceq

C = [9 (5.39 x 10-5 F)]

C = 4.85 x 10-4 F